OpenStudy (anonymous):

A man invests a certain amount of money at 2% interest and $800 more than that amount in another account at 4% interest. At the end of one year, he earned $92 in interest. How much money was invested in each account? $1,500 at 2%; $2,300 at 4% $1,400 at 2%; $2,200 at 4% $1,000 at 2%; $1,800 at 4%

OpenStudy (tkhunny):

Why not calculate the amounts and see? $1,500 at 2%; $2,300 at 4% ----- 2300 - 1500 = 800 Good ----- 1500*0.02 + 2300*0.04 = ?? $1,400 at 2%; $2,200 at 4% ----- 2200 - 1400 = 800 Good ----- 1400*0.02 + 2200*0.04 = ?? $1,000 at 2%; $1,800 at 4% ----- 1800 - 1000 = 800 Good ----- 1000/0.02 + 1800*0.04 = ??

OpenStudy (whpalmer4):

Not hard to write the equations to solve it: \[I = Prt\]First account:\[I_1 = Prt = P(0.02)(1)\]Second account:\[I_2 = (P+800)(0.04)(1)\]\[I_1+I_2 = 92\]\[0.02P + 0.04(P+800) = 92\]Solve for \(P\) for the amount in the first account, \(P+800\) is the amount in the second account.